# CRAFT: Training-Free Cascaded Retrieval for Tabular QA

### v1

@@ A.1 Stage-wise Results
Model R@1 R@10 R@50 R@100 R@1000 R@5000
Stage 1 34.38 72.90 91.62 94.89 98.91 99.56
     + Stage 2 36.65 82.91 96.08 97.82 98.91 -
        + Stage 3 41.13 87.16 96.84 - - -
CRAFT 41.13 87.16 96.84 - - -
Table 5: Table Retrieval Performance on NQ-Tables with various stages of CRAFT. R denotes Recall Score. Best performances are in Bold. ‘*’ denotes results reported for the full test-set with all 966 queries.
A.2 Ablation Study
Stage 1: Sparse Retrieval with Context Enrichment
We test the effect of table descriptions and query decomposition using BM25 and SPLADE. The results for Stage 1 are shown in Table 6 and Table 8. Both retrievers exhibit improvement when employing Question Decomposition. Specifically, SPLADE achieves gains of 2.18% and 4.03% at R@10 and R@50 respectively. However, incorporating Table Descriptions does not yield additional improvements for SPLADE, likely because the tables alone already exhaust the maximum context length of the SPLADE model, rendering extra descriptions redundant. Notably, SPLADE achieves a Recall@5000 of 99.59%, indicating performance saturation and leaving minimal scope for further enhancement. In contrast, BM25 demonstrates substantial improvement with the inclusion of table descriptions, clearly benefiting from context-enriched lexical matching.
Stage:2 Dense Semantic Reranking and Irrelevant Context Reduction
We apply all-mpnet-base-v2 to rerank the top 5000 tables using dense embeddings. Since dense embedding models have limited context length, we rank the rows of each table and retain the 5 most relevant rows per table. Table 9 show recall gains from filtering irrelevant rows, with R@10 rising from 79.21 to 82.91 and R@50 from 94.66 to 96.08. Dense scoring better captures semantic relevance missed in Stage 1.
Stage:3 Neural Reranking
Finally, we use text-embedding-3-small to rerank the top 300 and 100 tables. Table 10 shows the reranking results, Even without fine-tuning, neural reranking significantly sharpens final table rankings.
Table R@1 R@10 R@50 R@1000 R@5000
BM25
w/o Desc 17.95 34.38 47.44 70.73 82.48
+ Desc 22.74 45.16 62.57 86.61 93.14
SPLADE
w/o Desc 32.64 70.62 87.59 98.91 99.67
+ Desc 32.64 70.62 87.59 98.91 99.67
Table 6: Stage 1: Effect of Table Descriptions on Retrieval Performance over NQ-Tables. R denotes Recall Score. N1=169,898 Tables in Corpus. Best performances are in Bold.
Train Test
Total Query Count 9,594 966
Unique Query Count 9,594 919
Avg. # Words 8.94 8.90
Table Count 169,898 169,898
Avg. # Rows 10.70 10.70
Avg. # Columns 6.10 6.10
# Golden Tables / Query 1.00 1.05
Table 7: Statistics of the NQ-Tables benchmark. Test set contains 966 queries, out of which only 919 are unique.
Question R@1 R@10 R@50 R@1000 R@5000
BM25
Original 22.74 45.16 62.57 86.61 93.14
+ SubQues 24.15 53.54 71.38 91.08 96.84
SPLADE
Original 32.64 70.62 87.59 98.91 99.67
+ SubQues 34.38 72.90 91.62 98.91 99.56
Table 8: Stage 1: Effect of Query Decomposition- Retrieval Performance on NQ-Tables. R denotes Recall Score. N1=169,898 Tables in Corpus. Best performances are in Bold.
Table Text R@1 R@10 R@50 R@100 R@1000
Original* 37.64 79.76 94.45 97.82 99.45
     – Description 34.82 79.21 94.66 97.39 99.45
        – Irrelevant Rows 36.65 82.91 96.08 97.82 98.91
Table 9: Stage 2: Effect of Irrelevant Context Reduction- Retrieval using all-mpnet-base-v2 on NQ-Tables. Original* is title, header, all cell values, and description. N2=5000 Tables in Corpus. Best performances are in Bold.
N3 R@1 R@10 R@50
300 39.03 86.02 96.89
100 41.13 87.16 96.84
Table 10: Stage 3: Effect of Neural Reranking: Retrieval Performance using text-embedding-3-small model on NQ Tables. N3 is number of tables in corpus. Best performances are in Bold.
Configuration Parameter Value
Batch Size 144
Learning Rate $1\times 10^{-5}$
Training Epochs 50
Optimizer Adam
Initialization BERT-base
THYME Dropout ($p_{sem},p_{lex}$) 0.15
THYME FLOPS Reg. ($\lambda_{q},\lambda_{t}$) $1\times 10^{-4}$
Appendix B Further End-to-End Analysis
Retriever n=1 n=3 n=5 n=8 n=10
no_fewshot 39.65 39.29 33.33 46.04 44.06 37.70 45.09 45.98 41.47 43.78 44.49 49.99 42.58 44.12 50.69
fewshot_1 40.43 37.54 37.94 48.28 37.35 44.69 46.11 38.27 45.07 44.85 40.40 45.46 42.39 39.94 45.05
fewshot_2 41.56 42.73 38.98 49.29 37.77 43.16 46.69 36.83 46.02 45.35 40.54 45.03 43.53 39.35 44.86
fewshot_3 40.70 39.41 40.68 48.60 36.90 44.27 47.43 36.94 45.59 44.83 39.97 44.89 43.08 38.87 44.77
fewshot_4 41.58 40.96 40.99 48.77 37.41 43.86 47.59 36.72 46.19 45.56 40.55 44.64 43.72 40.06 43.37
fewshot_5 41.17 41.72 40.36 48.50 39.39 43.65 47.70 39.27 45.81 44.94 41.27 44.55 42.31 41.07 44.21
Table 12: F1 performance scores of different LLMs (Mistral-7B, Llama3-8B, and Qwen2.5-7B) across various retrieval configurations using CRAFT. Results are shown for increasing numbers of retrieved tables ($n=1,3,5,8,10$) and different few-shot prompting strategies. Each row represents a different few-shot variant, with no_fewshot indicating zero-shot performance.
End-to-End Performance Across Multiple Fewshot Examples.
To evaluate the scalability and robustness of CRAFT, we measure F1 performance of three pretrained LLMs, Mistral-7B, Llama3-8B, and Qwen2.5-7B, under zero-shot and few-shot prompting as the number of retrieved tables increases ($n=1,3,5,8,10$). As shown in Table 12, all models exhibit a clear upward trend in F1 score with deeper retrieval, demonstrating that structured multi-stage retrieval alone substantially narrows the gap with dataset-specific fine-tuning. However, incorporating a small number of in-context examples does not consistently yield significant improvements across models. Moreover, performance gains taper beyond $n=5$, suggesting an optimal retrieval depth that balances information richness with the risk of noise accumulation. This effect becomes more pronounced as both $n$ and the number of few-shot examples increase, leading to greater performance variability.
@@ Appendix B Further End-to-End Analysis
Token Consumption in Sub-Table vs. Full-Table Contexts.
@@
While deeper retrieval improves accuracy, computational efficiency remains critical for practical deployment. Table 13 compares the average token counts for each model when using only the top-$k$ sub-table context versus concatenating the entire table. Across all three models and $k\in{1,3,5}$, sub-table inputs reduce token length by over 70%. For example, Llama3 drops from 1782.80 tokens at $k=1$ to 362.70; Mistral from 3314.03 to 476.58; and Qwen2.5 from 1951.08 to 393.68. This substantial reduction directly lowers inference costs and improves runtime efficiency. By selectively presenting only the most row table fragments, CRAFT retains essential semantic content while avoiding the diminishing returns and latency overhead associated with full-table encoding.
B.1 Prompt Used For Experiments

### v2

@@ 4.4 Computational Efficiency
Computational Trade-offs .
Stage 2 employs a lightweight dense encoder to produce embeddings for the generated mini-tables on-the-fly during inference. While this increases per-query latency compared to retrieving from a pre-embedded corpus, it substantially reduces irrelevant content, thereby saving valuable context tokens for downstream processing. This design manages the trade-off between per-query cost and end-to-end performance by filtering noise early in the pipeline, which significantly benefits Stage 3 retrieval and LLM-based QA, both highly sensitive to token budgets. CRAFT’s modular design allows seamless integration of new embedding models, as efficiency is primarily governed by token budgets and per-query embedding calls.
End-to-End Efficiency.
While cascading retrieval improves performance with reduced offline computational overhead, computational efficiency also remains critical for the end-to-end QA task.
@@
Effect of Mini-Tables.
@@
Table 7 compares the average token counts for each model when using only the top-$k$ sub-table context versus concatenating the entire table. Across all three models and $k\in{1,3,5}$, sub-table inputs (CRAFT mini-tables) reduce token length by over 70%. For example, Llama3 drops from 1782.80 tokens at $k=1$ to 362.70; Mistral from 3,314.03 to 476.58; and Qwen2.5 from 1,951.08 to 393.68. This substantial reduction directly lowers inference costs and improves runtime efficiency. By selectively presenting only the most relevant table rows fragments, CRAFT retains essential semantic content while avoiding the diminishing returns and latency overhead associated with full-table encoding.
Effect of Token Reduction.
Figure 2 compares the end-to-end performance of three instruction-tuned models across varying context lengths on the NQ dataset. The curves reveal a clear trade-off between context breadth and noise accumulation: smaller retrieved sets (low n) often underperform because they may exclude the gold table, while larger retrieved sets (high n) increase the risk of including irrelevant content that dilutes attention and lowers F1. This pattern is particularly evident for Qwen2.5-7B-Instruct, where performance peaks at top_n = 9 using the Mini-Table (MT) configuration, showing that our multi-stage retrieval strategy effectively identifies an optimal context length that balances relevance and compactness. Stage 3 (MT) consistently achieves superior results at shorter context lengths, highlighting the benefit of reducing redundant information while retaining sufficient evidence for reasoning.
5 Ablation Study
[…]
NQ-Tables OTT-QA
Model R@1 R@10 R@50 R@100 R@500 R@5K R@1 R@10 R@50 R@100 R@500 R@5K
Stage 1 34.38 72.90 91.62 94.89 98.91 99.56 40.74 60.89 88.17 91.64 96.93 99.28
Stage 2 36.65 82.91 96.08 97.82 98.91 - 47.33 87.35 91.82 94.21 97.38 -
Stage 3 41.13 87.16 96.84 - - - 55.56 89.88 96.07 - - -
CRAFT 41.13 87.16 96.84 - - - 55.56 89.88 96.07
Table 8: Table Retrieval Performance on NQ-Tables and OTT-QA datasets with various stages of CRAFT. R denotes Recall Score. Best performances are in Bold. Boxed values indicate the number of tables passed to the next stage for re-ranking.
Error Analysis
To further understand the performance gaps, we provide a qualitative error analysis in Appendix D. We find that most failure cases arise from near-duplicate tables with highly similar schemas or ambiguous entity mentions, which are difficult to disambiguate without task-specific supervision. Additionally, even when the gold table is retrieved, using full tables can introduce noise that hinders answer extraction, whereas our mini-table representation isolates relevant rows and enables the LLM to generate correct answers more reliably.
Stage 1: Sparse Retrieval with Context Enrichment
We test the effect of table descriptions and query decomposition using BM25 and SPLADE. The results for Stage 1 are shown in Table 9 and Table 10. Both retrievers exhibit improvement when employing Question Decomposition. Specifically, SPLADE achieves gains of 2.18% and 4.03% at R@10 and R@50, respectively. With the table description, SPLADE achieved a consistent increase in recall. Notably, SPLADE achieves an R@5000 of 99.59%, indicating performance saturation and leaving minimal scope for further enhancement. In contrast, BM25 demonstrates substantial improvement with the inclusion of table descriptions, clearly benefiting from context-enriched lexical matching.
Table R@1 R@10 R@50 R@1k R@5k
BM25
w/o Desc 17.95 34.38 47.44 70.73 82.48
+ Desc 22.74 45.16 62.57 86.61 93.14
SPLADE
w/o Desc 31.06 69.05 86.85 98.45 99.59
+ Desc 32.64 70.62 87.59 98.91 99.67
Table 9: Stage 1: Effect of Table Descriptions on Retrieval Performance over NQ-Tables. R denotes Recall Score. N=169,898 Tables in Corpus.
Question R@1 R@10 R@50 R@1k R@5k
BM25
Original 22.74 45.16 62.57 86.61 93.14
+ SubQues 24.15 53.54 71.38 91.08 96.84
SPLADE
Original 32.64 70.62 87.59 98.91 99.67
+ SubQues 34.38 72.90 91.62 98.91 99.56
Table 10: Stage 1: Effect of Query Decomposition- Retrieval Performance on NQ-Tables. R denotes Recall Score. N=169,898 Tables in Corpus.
Stage 2: Dense Semantic Reranking and Irrelevant Context Reduction
We apply the Stage 2 dense model to rerank the top 5000 tables. Since dense embedding models have limited context length, we rank the rows of each table and retain the 5 most relevant rows per table. Table 11 shows recall gains from filtering irrelevant rows, with R@10 rising from 79.21 to 82.91 and R@50 from 94.66 to 96.08. Dense scoring better captures the semantic relevance missed in Stage 1.
Table Text R@1 R@10 R@50 R@100 R@1k
Original* 37.64 79.76 94.45 97.82 99.45
–Description 34.82 79.21 94.66 97.39 99.45
–Irrelevant Rows 36.65 82.91 96.08 97.82 98.91
Table 11: Stage 2: Effect of Irrelevant Context Reduction- Retrieval using NQ-Tables. Original* is the title, header, all cell values, and description. N2=5000 Tables in Corpus. Best performances are in Bold.
Stage 3: Neural Reranking
Finally, we rerank the top 300 and 100 mini-tables. Table 12 shows the reranking results. Even without fine-tuning, neural reranking significantly sharpens final table rankings.
N3 R@1 R@10 R@50
300 39.03 86.02 96.89
100 41.13 87.16 96.84
Table 12: Stage 3: Effect of Neural Reranking: Retrieval Performance using text-embedding-3-small model on NQ-Tables. N3 is the number of tables in the corpus. Best performances are in Bold.
These findings confirm the effectiveness of the cascaded design, which begins with broad lexical and semantic coverage, progressively refines relevance through dense pruning, and culminates in lightweight neural reranking for precise top-rank. Collectively, they highlight how each stage contributes a distinct function within the pipeline, yielding an accurate and computationally efficient retrieval framework.
6 Conclusion
We present CRAFT, a modular, training-free cascaded retrieval framework for Table Question Answering that achieves competitive performance compared to state-of-the-art systems that require dataset-specific fine-tuning. CRAFT achieves new state-of-the-art retrieval Recall@1 on NQ-Tables, maintains high robustness against query paraphrasing, and delivers strong end-to-end QA performance when paired with modern LLMs. By decoupling retrieval from domain-specific fine-tuning, CRAFT offers a scalable and adaptable paradigm for dynamic tabular domains where rapid deployment and data scarcity are critical issues. Future work will extend CRAFT’s capabilities to enhance its reasoning through a fully integrated, multi-hop retrieval system to tackle complex benchmarks, such as OTT-QA, without relying on oracle passages. These directions will further advance CRAFT toward a unified, efficient, and generalizable solution for open-domain table understanding.
7 Limitations
While this study demonstrates the effectiveness of a training-free, cascaded approach, its scope has certain boundaries. The evaluation centers on single-hop question answering over NQ-Tables and OTT-QA, leaving multi-hop reasoning tasks that demand synthesis across multiple tables and passages unexplored. One promising extension adds a lightweight contrastive objective after Stage 3 to sharpen R@1 by separating structurally similar yet semantically distinct tables, which we leave for future work.
The end-to-end QA evaluation for OTT-QA ran in an oracle setting, using gold passages to isolate retrieval performance, falling short of a full open-retrieval pipeline. Our approach targets text-based table retrieval and does not handle multimodal tables containing images or other non-textual data. Finally, although CRAFT’s offline pre-computation runs efficiently, the pipeline’s multi-stage design pushes per-query processing time upward.
8 Acknowledgements
We thank the anonymous reviewers and the Area Chair for their constructive feedback. We are grateful to the members of the Complex Data Analysis and Reasoning Lab for their helpful discussions and contributions to the project’s ideation. We also thank Aman Singh and Prasham Titiya for their insightful input during the early stages of this work. We also acknowledge the support of the USA National Science Foundation under grant #2047488 and the Rensselaer-IBM Future of Computing Research Collaboration (FCRC).
9 Ethics Statement
This research adheres to standard ethical practices in NLP. We built our approach using publicly available datasets, NQ-Tables and OTT-QA, that contain no sensitive or personally identifiable information. For generating table titles, descriptions, and query expansions, we leveraged Gemini 1.5 Flash alongside other embedding models and language models, all used within their specified usage guidelines. Since our framework is training-free, we didn’t need to collect or annotate new data. That said, we recognize that our method can reflect biases present in the underlying datasets and pre-trained models. To address this, we kept our experiments in controlled settings and report results with full transparency. We’ve included enough implementation details to make our work reproducible, and any AI-assisted writing tools were used solely to refine clarity and all outputs were carefully reviewed and verified by the authors.
Appendix A Hyperparameter Configuration
Configuration Parameter Value
Batch Size 144
Learning Rate $1\times 10^{-5}$
Training Epochs 50
Optimizer Adam
Initialization BERT-base
THYME Dropout ($p_{sem},p_{lex}$) 0.15
THYME FLOPS Reg. ($\lambda_{q},\lambda_{t}$) $1\times 10^{-4}$
Table 13 outlines the training and inference configuration adopted for baseline retrievers. These configurations were chosen to ensure stable convergence across both datasets while maintaining generalization in retrieval performance. The combination of moderate dropout, lightweight regularization, and a conservative learning rate enabled consistent optimization across heterogeneous data distributions, striking a balance between efficiency and representational expressiveness. For Re2G , we directly used the publicly released checkpoints fine-tuned on NQ Dataset.
Appendix B Prompt Used For Experiments
@@
Appendix C Further End-to-End Analysis
Retriever n=1 n=3 n=5 n=8 n=10
Zero fewshot 39.13 38.31 39.73 45.28 40.76 43.52 44.53 40.55 46.49 44.01 41.11 46.53 41.50 39.15 47.24
1-Fewshot 40.43 37.54 37.94 48.28 37.35 44.69 46.11 38.27 45.07 44.85 40.40 45.46 42.39 39.94 45.05
2-Fewshot 41.56 42.73 38.98 49.29 37.77 43.16 46.69 36.83 46.02 45.35 40.54 45.03 43.53 39.35 44.86
3-Fewshot 40.70 39.41 40.68 48.60 36.90 44.27 47.43 36.94 45.59 44.83 39.97 44.89 43.08 38.87 44.77
4-Fewshot 41.58 40.96 40.99 48.77 37.41 43.86 47.59 36.72 46.19 45.56 40.55 44.64 43.72 40.06 43.37
5-Fewshot 41.17 41.72 40.36 48.50 39.39 43.65 47.70 39.27 45.81 44.94 41.27 44.55 42.31 41.07 44.21
Table 14: F1 performance scores of different LLMs (Mistral-7B, Llama3-8B, and Qwen2.5-7B) across various retrieval configurations using CRAFT. Results are shown for increasing numbers of retrieved tables ($n=1,3,5,8,10$) and different few-shot prompting strategies. Each row represents a different few-shot variant, with zero_fewshot indicating zero-shot performance.
Retriever n=1 n=3 n=5 n=8 n=10
Mistral Llama3 Qwen2.5 Mistral Llama3 Qwen Mistral Llama3 Qwen Mistral Llama3 Qwen Mistral Llama3 Qwen
Small Models
Gold 43.60 41.10 52.87 - - - - - - - - - - - -
Stage 2(MT) 38.83 35.58 46.49 41.71 25.78 48.37 40.54 29.11 48.97 38.09 30.38 47.76 36.96 29.53 46.67
Stage 2(FT) 37.72 34.62 46.69 39.46 25.73 47.31 38.40 29.09 49.06 36.20 29.53 47.32 35.12 28.35 46.02
Stage 3 Gemini(MT) 39.38 36.60 49.38 41.02 26.99 50.15 39.14 29.81 48.88 36.16 31.59 46.77 35.44 30.75 47.93
Stage 3 Gemini(FT) 38.28 36.08 48.54 39.38 26.29 49.91 38.10 29.24 49.16 34.95 30.00 47.16 34.32 30.03 46.93
Large Models
Gold 61.38 59.30 65.31 - - - - - - - - - - - -
Stage 2(MT) 54.47 52.06 56.36 56.90 50.52 61.59 57.18 50.44 62.22 54.84 51.43 62.03 53.11 51.45 61.34
Stage 2(FT) 54.80 51.50 56.79 57.52 50.52 60.74 58.12 50.94 62.59 55.66 51.55 61.45 53.25 51.00 61.69
Stage 3 Gemini(MT) 55.43 53.24 59.24 59.41 51.95 63.55 56.77 50.99 63.34 54.12 51.42 63.16 52.86 50.61 62.80
Stage 3 Gemini(FT) 55.06 53.28 58.35 59.21 51.17 62.95 57.27 51.13 62.17 54.65 51.08 61.92 53.72 50.69 60.06
Table 15: F1 performance scores of different LLMs (Top(Small Models): Mistral-7B-Instruct, Llama3-8B-Instruct, and Qwen2.5-7B-Instruct, and Bottom(Large Models): Mistral-Small-Instruct, Llama-3.3-70B-Instruct, and Qwen2.5-72B-Instruct) across various retrieval configurations using CRAFT for OTT-QA dataset. Results are shown for increasing numbers of retrieved tables ($n=1,3,5,8,10$) and different table retrievals. Each row represents a different table retrieved using different table embeddings, with Gold indicating end-to-end score with the gold table and passage.
Figure 6: End-to-end performance comparison across context lengths with OTTQA dataset. The figure illustrates the relationship between average context length and F1 score for three models(small size models), evaluated over 10 different $n$-retrieved tables with the OTT-QA datasets.
Figure 7: End-to-end performance comparison across context lengths with OTT-QA dataset. The figure illustrates the relationship between average context length and F1 score for three models(large size models), evaluated over 10 different $n$-retrieved tables with the OTT-QA datasets.
End-to-End OTT-QA Performance
Table 15 presents the end-to-end F1 performance of different instruction-tuned LLMs on the OTT-QA dataset using our multi-stage retrieval framework. It compares multiple retrieval strategies across varying depths, highlighting how retrieval design impacts downstream QA accuracy.
We consider the following retrieval settings:
  • Gold: Uses the gold table (oracle evaluation setup) paired with gold passages to isolate the effect of table retrieval and provide an upper-bound reference.
  • FT (Full Table): Uses the top-$n$ retrieved table IDs (from Stage 2 or Stage 3) in their full form as context for answering the question.
  • MT (Mini Table): Uses compact representations of the retrieved tables (based on table IDs), retaining only top(k=5) rows to reduce noise and improve reasoning efficiency.
Since OTT-QA is a complex hybrid multi-hop benchmark that requires reasoning across both tables and passages, we employ an oracle evaluation setup that pairs the retrieved tables with the corresponding gold passages.
Across both small and large models, highlighted by Figure 6 and Figure 7 respectively, performance trends reveal a clear trade-off between recall and precision. For smaller $n$, the retrieved tables may not include the gold table, leading to suboptimal results. Conversely, higher $n$ increases the likelihood of retrieving the relevant table but also introduces redundant or noisy tables that dilute reasoning accuracy. Among all settings, Stage 3 (MT) consistently achieves strong performance, particularly for Qwen2.5-7B-Instruct, which reaches its best score at n = 3–5, closely matching the gold-table upper bound. This demonstrates the benefit of the Mini-Table (MT) strategy, which compactly represents relevant content while minimizing contextual noise. For large-scale models such as Qwen2.5-72B-Instruct, Stage 3 Gemini (MT) yields the most balanced results, highlighting how our approach generalizes effectively across model sizes by achieving a “sweet spot” between retrieval breadth and contextual precision.
Appendix D Error Analysis
Our error analysis reveals several consistent patterns in CRAFT’s behavior under imperfect retrieval conditions. We find that retrieved tables often exhibit high topical similarity to the query, yet do not necessarily contain the specific evidence required for multi-hop reasoning, indicating that surface-level relevance is an unreliable proxy for answerability. Moreover, the rank of the gold table does not directly correlate with end-to-end QA performance, therefore even lower-ranked tables can support correct predictions when paired with effective mini-table representations that isolate relevant evidence. We also observe that standard retrieval metrics such as R@1 and R@5 tend to underestimate overall system capability, as downstream reasoning components can compensate for imperfect retrieval by extracting useful signals from non-top-ranked candidates. Finally, using full tables as input introduces substantial noise, frequently leading to missed answers or spurious predictions, whereas mini-table representations reduce irrelevant context and enable more accurate reasoning. Taken together, these findings suggest that improving QA performance requires joint optimization of retrieval and reasoning, with intermediate representations such as mini-tables playing a central role.
Case Study I: Full Table Fails Despite Correct Retrieval.
Table 17 presents an example where the gold table (1997_AFL_draft_1) is ranked third among the top-5 retrieved tables, yet the full-table approach fails to extract the correct answer, responding with Pass (i.e., unable to find relevant information). In contrast, the mini-table design successfully identifies Jamie Elliott as the answer. This illustrates how CRAFT can exhibit low Recall@1 while still producing correct final answers through focused context construction.
Attribute Value
QID e4197b9b9bd82bdd
Question Who was the St. Kilda pick 11
in the 1997 AFL draft?
Gold Answer Jamie Elliott
Gold Table 1997_AFL_draft_1
Gold Table Rank 3
Top 5 Retrieved TableIDs
1 1998_AFL_Draft_3
2 List_of_St_Kilda_Football_Club_
records_and_statistics_8
3 1997_AFL_draft_1 ✓
4 1998_AFL_Draft_2
5 List_of_St_Kilda_Football_Club_
records_and_statistics_9
LLM QA Step
Full Table Pass (no answer found) $\times$
Mini-Table Jamie Elliott ✓
Table 16: Case Study I: The gold table is ranked third. The full-table approach fails to extract the correct answer, while the mini-table design succeeds by isolating the relevant row.
Attribute Value
QID bde8169851a494cf
Question What album did Elisa’s most
popular original song come out on?
Gold Answer Asile’s World
Gold Table List_of_songs_recorded_by_Elisa_0
Gold Table Rank 2
Top-5 Retrieved Table Ids
1 List_of_songs_recorded_by_Elisa_1
2 List_of_songs_recorded_by_Elisa_0  ✓
3 Elisa_(Italian_singer)_1
4 Elisa_(Italian_singer)_0
5 Raquel_Olmedo_1
LLM QA Step
Full Table Lotus (2003) $\times$
Mini-Table Asile’s World (2000) ✓
Table 17: Case Study II: Even with the gold table ranked second, the full-table approach hallucinates an incorrect answer, while the mini-table approach correctly isolates the relevant information.
Case Study II: Full Table Hallucination with Closely-ranked Retrieval.
Table 17 presents a more nuanced failure mode. The gold table (List_of_songs_recorded_by_Elisa_0) is ranked second, which is a relatively strong retrieval result, yet answer quality diverges sharply between the two approaches. The full-table approach returns Lotus (2003), a plausible but incorrect answer drawn from a different row in the retrieved context. The mini-table approach correctly returns Asile’s World (2000) by isolating the relevant row corresponding to Elisa’s most popular original song.
Taken together, these case studies highlight that the benefit of mini-tables is not limited to weak retrieval scenarios. Even when relevant tables are retrieved at high ranks, full-table context can introduce sufficient noise through topically similar but irrelevant rows to cause model failure or hallucination. Mini-table construction acts as a targeted denoising step, enabling the downstream QA model to focus on the precise evidence needed for multi-hop inference.